# Diffusion-Reward Adversarial Imitation Learning

### v1

@@ Appendix D Converged performance
Appendix E Model architecture
This section presents the model architecture of all the experiments. Section E.1 describe the model architecture of all methods used in Section 5.3.
Table 2: Model architectures of policies and discriminators. We report the architectures used for all the methods on all the tasks. Note that $\pi$ denotes the neural network policy, $D$ represents a multilayer perceptron discriminator used in GAIL and WAIL, and $D_{\phi}$ represents a diffusion model discriminator used in DiffAIL and our method DRAIL.
@@ Appendix E Model architecture
E.1 Model architecture of DRAIL, DiffAIL, and the baselines
In Section 5.3, we conducted a comparative analysis between our proposed DRAIL, along with several baseline approaches (BC, Diffusion Policy, GAIL, WAIL, and DiffAIL) across six diverse environments. We applied Multilayer Perceptron (MLP) for the policy of BC, the conditional diffusion model in Diffusion Policy, as well as the policy and the discriminator of GAIL and WAIL. For DiffAIL and our proposed DRAIL, MLPs were employed in the policy and diffusion model of the diffusion discriminative classifier. The activation functions used for the MLPs in the diffusion model were ReLU, while hyperbolic tangent was employed for the others. The total timestep $T$ for all diffusion models in this paper is set to 1000 and the scheduler used for diffusion models is cosine scheduler [35]. Further details regarding the parameters for the model architecture can be found in Table 2.
@@ E.1 Model architecture of DRAIL, DiffAIL, and the baselines
GAIL.
The detailed model architecture for GAIL is presented in Table 2.
@@
E.2 Image-based model architecture of DRAIL, DiffAIL, and the baselines
@@
Discriminator of GAIL & WAIL.
The discriminators of GAIL and WAIL are similar to the policy model; the only difference is that the last linear layer outputs a 1-dimensional value that indicates the probability of a given state-action pair being from the expert demonstrations.
@@
Appendix F Training details
F.1 Training hyperparamters
@@
F.2 Reward function details

### v4

@@ Appendix E Converged performance
Appendix F Model architecture
This section presents the model architecture of all the experiments. Section F.1 describe the model architecture of all methods used in Section 5.3.
Table 2: Model architectures of policies and discriminators. We report the architectures used for all the methods on all the tasks. Note that $\pi$ denotes the neural network policy, $D$ represents a multilayer perceptron discriminator used in GAIL, GAIL-GP, and WAIL, and $D_{\phi}$ represents a diffusion model discriminator used in DiffAIL and our method DRAIL.
@@ Appendix F Model architecture
F.1 Model architecture of DRAIL, DiffAIL, and the baselines
In Section 5.3, we conducted a comparative analysis between our proposed DRAIL, along with several baseline approaches (BC, Diffusion Policy, GAIL, GAIL-GP, WAIL, and DiffAIL) across six diverse environments. We applied Multilayer Perceptron (MLP) for the policy of BC, the conditional diffusion model in Diffusion Policy, as well as the policy and the discriminator of GAIL, GAIL-GP, and WAIL. For DiffAIL and our proposed DRAIL, MLPs were employed in the policy and diffusion model of the diffusion discriminative classifier. The activation functions used for the MLPs in the diffusion model were ReLU, while hyperbolic tangent was employed for the others. The total timestep $T$ for all diffusion models in this paper is set to 1000 and the scheduler used for diffusion models is cosine scheduler [38]. Further details regarding the parameters for the model architecture can be found in Table 2.
@@ F.1 Model architecture of DRAIL, DiffAIL, and the baselines
GAIL & GAIL-GP.
The detailed model architecture for GAIL and GAIL-GP is presented in Table 2. For GAIL-GP, the gradient penalty is set to 1 across all environments.
@@
F.2 Image-based model architecture of DRAIL, DiffAIL, and the baselines
@@
Discriminator of GAIL, GAIL-GP, & WAIL.
The discriminators of GAIL, GAIL-GP, and WAIL are similar to the policy model; the only difference is that the last linear layer outputs a 1-dimensional value that indicates the probability of a given state-action pair being from the expert demonstrations.
@@
Appendix G Training details
G.1 Training hyperparamters
@@
G.2 Reward function details